OpenStudy (anonymous):

AP Calculus: Slope Fields/ Separate Variables Find the equation of the curve, given the derivative and the indicated point on the curve. dy/dx=lnx/x

OpenStudy (anonymous):

What do you find confusing about how to solve the differential equation? Have you already learned the method of separation of variables?

OpenStudy (anonymous):

kinda but i was absent that day when we learned it in class :\

OpenStudy (anonymous):

The first step in any differential equation problem is to determine whether you can separate variables. That is to say, can you move all the terms with "dy" and "y" to one side, and "dx" and "x" to the other?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Now do so; in this case, you can treat "dx" as its own term and multiply both sides by it. Then you've got\[dy = \frac{ \ln{x} dx }{ x }\] Now integrate both sides and you're virtually done!

OpenStudy (anonymous):

so far i have \[\int\limits_{}^{} dy = \int\limits_{}^{} \frac{ lnx }{ x }dx\]

OpenStudy (anonymous):

i dont know how to integrate \[\int\limits_{}^{} \frac{ lnx }{ x } dx\] though :\

OpenStudy (anonymous):

integration by parts?

OpenStudy (anonymous):

would i use substituion?

OpenStudy (anonymous):

There's no harm in trying a few different methods! This one's simple enough that it shouldn't take too much time.

OpenStudy (anonymous):

so would i use the product rule?

OpenStudy (anonymous):

I think that'd be a bit tedious, as integration by parts problems usually are. Try a substitution.

OpenStudy (anonymous):

so then i use substitution for this?

OpenStudy (anonymous):

so you split it up to\[\int\limits_{}^{} \ln x \times \frac{ 1 }{ x } dx\]

OpenStudy (anonymous):

It's a good rule of thumb to try substitution when you see fractions in the integrand. The problem simplifies very quickly with a particular substitution.

OpenStudy (anonymous):

ok so when i get \[\int\limits_{}^{} \ln x \times \frac{ 1 }{ x } dx \] what do i next to integrate it?

OpenStudy (anonymous):

Have you tried a substitution? Think of the most obvious one that might work.

OpenStudy (anonymous):

yeah i have but idk how to get x

OpenStudy (anonymous):

so let \[u=\ln x\] right?

OpenStudy (anonymous):

You tell me if it's right.

OpenStudy (anonymous):

Carry it out. :)

OpenStudy (anonymous):

ok but how would i find x

OpenStudy (anonymous):

is x \[\frac{ u }{ \ln }= x\]

OpenStudy (anonymous):

The goal of substitution is to eliminate your original variable, in this case it was x. Differentiate your substitution to get \[du = \frac{1}{x} dx\] and you can rearrange it to get dx. Then, put that dx into your original integral and every term with x in it will have disappeared and you will have a simple integral in u and du.

OpenStudy (anonymous):

so u isn't \[u=\ln x\]

OpenStudy (anonymous):

It is, but you have to differentiate that equation you've made.

OpenStudy (anonymous):

but what would i substitue for x?

OpenStudy (anonymous):

Once you differentiate your substitution, you get the statement I said above, and you acquire \[dx = x du\] which you can put into your original integral. The x's will cancel out and your ln(x) turns into a u, as you already said it would.

OpenStudy (anonymous):

You end up with \[\int\limits u \ du\] which I think you know how to handle.

OpenStudy (anonymous):

i'm still confused

OpenStudy (anonymous):

for substitution u have to find u, x, and du. \[u = \ln x\]

OpenStudy (anonymous):

Yes. Differentiate it, and what do you have?

OpenStudy (anonymous):

\[u=lnx\] would you divide ln to get x by itself?

OpenStudy (anonymous):

No, ln(x) is a function of x. You can't divide by ln. What is the derivative of ln(x)?

OpenStudy (anonymous):

\[\frac{ 1 }{ x }\]

OpenStudy (anonymous):

So you have du/dx = 1/x. Now you want to find what dx is, in order to replace it in your original integral.

OpenStudy (anonymous):

so i divide by du so i would get \[\frac{ 1 }{ dx }= \frac{ 1 }{ x \times du }\]

OpenStudy (anonymous):

Yes, and you can flip both sides to get dx = x*du. Now, you have dx = x*du, and you have u = ln(x). Put those both into your integral from before.

OpenStudy (anonymous):

are you sure substitution is the only way for this problem?

OpenStudy (anonymous):

cause i don't really like using substituion. My Calculus teacher said that should be our last resolution in dealing with integrals..

OpenStudy (anonymous):

I'm sure it's not the only way, but it certainly is the fastest and simplest in most cases.

OpenStudy (anonymous):

when you explain your way of method of substitution you confuse me :\

OpenStudy (anonymous):

when we do substitution the equation usually has a square root

OpenStudy (anonymous):

can you show me another method on how to get the integral of \[\int\limits_{}^{} \frac{ lnx }{ x }dx\] cause i'm pretty sure my teacher doesn't like seeing us do substituion with these types of problems unless it involves a square root in them.

OpenStudy (anonymous):

Your teacher shouldn't have a preference either way, it's just as valid a method as any other, and you're less likely to make mistakes with this substitution because integration by parts tends to be easy to lose track of.

OpenStudy (anonymous):

well can you show me the product rule for this one or anything besides substitution? I have a test on this tomorrow :\

OpenStudy (anonymous):

Using product rule, choosing u = ln(x) and dv/dx = 1/x, we have \[\int\limits \ln(x) \frac{ 1 }{ x } dx = \ln(x) \ln(x) - \int\limits \ln(x) \frac{1}{x} dx\] Then, add the integral on the right-hand side to the left hand side, and divide the whole expression by two to get\[y = \frac{\ln^2(x)}{2} + c\]

OpenStudy (anonymous):

A great rule I was taught to determine which term to set equal to u is "LIPET" - Logarithms, inverse-trigonometric functions, polynomials, exponentials, and trigonometric functions. Whichever comes first in the lineup will be the one you should pick for your "u", and it will make your integral easier to solve.

OpenStudy (anonymous):

yeah i still dont get it haha :P